1* The Retread Tire Company recaps tires. The fixed annual cost of the recapping opration is $60,000
The variable cost of recapping a tire is $9.
The company charges $25 to recap a tire.
a. For an annual volume of 12,000 tires, determine the total cost, total revenue, and profit.
b. Deternine the annual break-even volume for the Retead Tire Company operation.
2* If the Retread Tire Company in the above problem changes its pricing for recapping a tire from $25 to $31, what effect will the change have on the break-even volume?

Answer by richwmiller(17219)

1a)ar=25*t
ar=25*12k=300k
ac=9*t
ac=9*12k+60k=168k
ap=ar-ac
ap=25*12k-9*12k-60k
ap=300k-168k
ap=132l
1b)
25t=60000+9*t
16t=60000
be t=3750
2)31t=60000+9*t
22t=60000
t=2727
The volume needed to break even goes down when the selling price goes up to $31